OpenStudy (tylermckinney16):

Help !! Identify the vertex, focus, and directrix of this parabola: y=2(x-3)^2-4.

OpenStudy (tylermckinney16):

@Directrix

Directrix (directrix):

The equation is given in vertex form so that you can pick out the coordinates of the vertex of the parabola. y = 2*(x-3) ^2 -4 Vertex Form of Parabola y = a * (x - h)^2 + k where (h.k) are the coordinates of the vertex of the parabola.

OpenStudy (tylermckinney16):

Ok

OpenStudy (tylermckinney16):

That the answer or is there anything else?

Directrix (directrix):

You have to get the coordinates of the vertex. Then, the coordinates of the Focus, and the equation of the Directrix.